A manufacturer has determined that the revenue received from selling x items of a product is given by R(x) = -0.3x2 + 200x and the cost to produce x items of the product is given by C(x) = 80x + 9400. Assuming that all of the products produced can be sold, how many should be produced to maximize the profit? (Hint: profit = revenue - cost)
I have such a hard time with word problems. I reall appreciate the help

Answer by scott8148(5895)

let P(x) equal profit ... P(x)=R(x)-C(x) ... P(x)=-0.3x^2+200x-(80x+9400) ... P(x)=-0.3x^2+120x-9400

the maximum lies on the axis of symmetry of the graph and is defined by x=-b/2a

in this case, x=-120/(2(-0.3)) ... x=200